Because elderly people may have difficulty standing to have their height measured, a study looked at the relationship between overall height and height to the knee. Here are data (in centimeters) for five elderly men:

Knee Height xx 57.7 46.7 42.7 47.2 57.6
Height yy 192.9 154.6 146.7 162.4 168.2


What is the equation of the least-squares regression line for predicting height from knee height?

ANSWER: y^=
Answer #1

Sum of X = 251.9
Sum of Y = 824.8
Mean X = 50.38
Mean Y = 164.96
Sum of squares (SSX) = 188.348
Sum of products (SP) = 414.416

Regression Equation = ŷ = bX + a

b = SP/SSX = 414.42/188.35 = 2.2003

a = MY - bMX = 164.96 - (2.2*50.38) = 54.1105

Ans is ŷ = 2.2003X + 54.1105